NvidiaIntely (nvidiaintely):

Describe the different ways to shift the equilibrium to the left for the following reaction: 2 A (s) + 2 BC (aq) arrow 2 AB (aq) + C2 (g). Justify your answer. ΔH = −352 kJ

NvidiaIntely (nvidiaintely):

@sweetburger @Kevin @jhonyy9

NvidiaIntely (nvidiaintely):

@Nnesha

NvidiaIntely (nvidiaintely):

@sshayer @AloneS

OpenStudy (kevin):

Sorry for waiting, I had a job to do before..

OpenStudy (kevin):

2 A (s) + 2 BC (aq) arrow 2 AB (aq) + C2 (g). ΔH = −352 kJ This is exothermic reaction, because ΔH < 0 The different ways to shift the equilibrium to the left for the following reaction: 1. "decrease reactant concentration, that's AB and C2 concentration" OR "increase product concentration, that's A and BC concentration" 2. exothermic : AB and C2 endothermic : A and BC So to shift equilibrium to left we need to make the equilibrium to shift to endothermic by : "Increase the temperature" 3. 2 A (s) + 2 BC (aq) arrow 2 AB (aq) + C2 (g) reactant coefficient : 2 + 2 = 4 product coefficient : 2 + 1 = 3 If we "decrease the pressure or increase the volume", then the equilibrium will shift to reaction with highest coefficient.

NvidiaIntely (nvidiaintely):

THanks Kevin!

OpenStudy (kevin):

You're welcome :D